# A Numerical Study on Gauge Symmetry of Electroweak Amplitudes

### v2

@@ Abstract
Electroweak (EW) amplitudes in the gauge-Goldstone 5-component formalism have a distinctive property: gauge symmetry is imprinted in the amplitudes, manifested as the massive Ward identity (MWI) $k^{M}\mathcal{M}_{M}=0$. In this paper, we use the HELAS package to numerically study gauge symmetry in EW amplitudes. First, we directly test gauge symmetry by examining the MWI of amplitudes. Second, we modify the couplings within a vertex and of vertices to check if and how the MWI changes. Third, we test gauge symmetry by considering the couplings modified by operators from the standard model effective field theory (SMEFT). Similar to the standard model, there are relations between different couplings that are protected by gauge symmetry. We find that if we modify the couplings to deviate from the relations, the MWI is violated. On the other hand, the MWI is restored when the relations between couplings reduce to those in the SMEFT.
@@ I Introduction
The gauge cancellation of electroweak (EW) scattering amplitudes in the standard model (SM) Glashow:1961tr ; Weinberg:1967tq ; Salam:1968rm has been both a theoretical problem hindering physical analysis and a practical problem in numerical calculations. In recent years, a new framework of computing EW amplitudes based on Goldstone equivalence was proposed and implemented in the software package HELAS Chen:2022gxv . The main ingredients of this framework are: 1) taking Goldstone equivalence Cornwall:1974km ; Chanowitz:1985hj ; Gounaris:1986cr , so that the polarization vector of a massive vector boson has no $k^{\mu}$ term, instead is composed of its Goldstone component and a remnant gauge term; 2) combining the gauge components and their Goldstone component of fields or polarization vectors into single 5-component objects at the level of polarization vectors, propagators, and vertices; 3) imposing a special light-cone gauge defined by the gauge direction $n^{\mu}=(1,-{\vec{k}}/{|\vec{k}|})$, dubbed the Feynman diagram gauge Chen:2022gxv or Goldstone equivalence gauge Chen:2016wkt . It has been demonstrated in Ref. Chen:2022gxv with many examples that EW amplitudes in this new scheme do not have gauge cancellation, thereby solving a long-standing problem. Further studies on this topic can be found in Refs. Hagiwara:2024ron ; Chen:2022xlg ; Furusato:2024uli ; Hagiwara:2024xdh ; Zheng:2024lwl ; Furusato:2024uli ; Furusato:2024ghr ; Jeong:2025hwj . Earlier but incomplete treatment of the 5-component formalism can be found in Kunszt:1987tk ; Wulzer:2013mza ; Cuomo:2019siu .
@@
The new framework in Ref. Chen:2022gxv is essentially a reorganization and combination of “pieces” from existing Feynman rules. Although ideal to be combined with the Feynman diagram gauge, it can also be applied to other gauges, such as the Feynman gauge and others. Therefore, it is necessary to distinguish the 5-component framework from any specific choice of gauge. Additionally, since the framework relies on the Goldstone equivalence theorem Cornwall:1974km ; Chanowitz:1985hj ; Gounaris:1986cr in an essential way, it is appropriate to refer to it as the Goldstone equivalence (GE) representation of Feynman rules, as we will do from now on. Correspondingly, the standard Feynman rules will be called the gauge representation, as they exclusively treat the physical content of massive vector bosons as quanta of gauge fields.
@@
Although the absence of gauge cancellation is crucial for practical applications, it is not the only important property of the GE representation of EW Feynman rules. Another intriguing property is that the GE representation directly imprints gauge symmetry at the level of amplitudes, as expressed by the massive Ward identity (MWI), which is named in analogue to its massless counterpart. Specifically, for a scattering amplitude $\varepsilon^{\mu(*)}(k)\mathcal{M}_{\mu}$ involving an external vector boson $V$ with mass $m_{V}$ and 4-momentum $k^{\mu}$, the MWI reads Cornwall:1974km ; Chanowitz:1985hj ; Gounaris:1986cr ; Bagger:1989fc
@@
where the -- ($+$) sign corresponds to the case that the vector boson is in the initial (final) state, and $\mathcal{M}(\varphi)$ represents the amplitude obtained by replacing the vector boson with the corresponding Goldstone boson $\varphi$. The MWI (1) is a fundamental identity for a spontaneously broken gauge theory, e.g., the EW gauge theory. The main goal of this work is to study gauge symmetry of EW amplitudes using the numerical tool HELAS Hagiwara:1990dw ; Murayama:1992gi ; Chen:2022gxv .
Similar to the Ward identity in gauge theories with massless gauge bosons, the MWI guarantees precise relations among different diagrams of an amplitude, which further imply exact relations among different vertices, or equivalently, different couplings within the same theory. Taking the example of $W^{+}W^{-}\rightarrow W^{+}W^{-}$ with the helicity combination of $\mathrm{TTTT}$, where $\mathrm{T}$ represents a transverse polarization, the amplitude includes three vertices $WWZ$, $WWA$, and $WWWW$ with couplings $g_{WWZ}$, $g_{WWA}$, and $g_{WWWW}$. The Ward identity requires
@@
Conversely, any deviation from this relation would result in violations of the gauge symmetry and the Ward identity. When the helicity combination involves longitudinal components, the precise relations the MWI brings about not only involve couplings of gauge bosons, but also those of Goldstone bosons and the Higgs boson. Thus, the structure of gauge symmetry is much richer in massive amplitudes, on which we will devote much space to study. Moreover, the gauge symmetry and the MWI are not limited to the SM but can also be applied to other theories. One theory of particular interests is the standard model effective field theory (SMEFT) Coleman:1969sm ; Callan:1969sn ; Weinberg:1980wa ; Leung:1984ni ; Buchmuller:1985jz ; Grzadkowski:2010es , which is a direct extension of the SM.
The remaining of the paper is organized in the following way. In Section II, we provide a brief introduction to the GE representation of EW interactions, emphasizing on the relations among different couplings and parameters. In Section III, we directly test the EW gauge symmetry with the MWI (1) across multiple processes and helicity combinations. In Section VI, we examine the EW gauge symmetry through anomalous couplings, including both within individual vertices and among multiple vertices. Finally in Section V, we study the connection between anomalous couplings (Higgs self-couplings and Yukawa couplings) and certain SMEFT operators using the MWI. We will demonstrate that when the anomalous couplings are adjusted according to the SMEFT operators, the MWI is satisfied, ensuring gauge invariance and self-consistency of the theory.
@@ II GE representation of EW interactions
II.1 Polarization Vectors in the 5-component Formalism
The central identity for our program is the massive Ward identity (1), which essentially states the equivalence between the $k^{\mu}$ terms in gauge fields and the corresponding Goldstone fields. Moreover, the longitudinal polarization vector of a massive vector boson $V$ with 4-momentum $k^{\mu}$ can be decomposed into
@@ II.1 Polarization Vectors in the 5-component Formalism
with $n^{\mu}=(1,-{\vec{k}}/{|\vec{k}|})$ and the subscript $\mathrm{L}$ representing a longitudinal polarization. We can eliminate the $k^{\mu}$ term in $\epsilon_{\mathrm{L}}^{\mu}(k)$ by using the MWI, resulting in the amplitude with a longitudinal vector boson to become
@@
The generalization to multiple vector bosons is straightforward, but it leads to an excessive number of terms, making calculations cumbersome. To address this issue, we define 5-component longitudinal polarization vectors in the GE representation by combining the gauge boson wave function and the Goldstone boson “wave function” as
$\epsilon^{M}_{\mathrm{L}}(k)$ $\equiv\left(-\frac{m_{V}n^{\mu}}{n\cdot k},\leavevmode\nobreak\ i\right)\ \ \ \text{for initial state,}$, (5)
$\epsilon^{*M}_{\mathrm{L}}(k)$ $\equiv\left(-\frac{m_{V}n^{\mu}}{n\cdot k},\leavevmode\nobreak\ -i\right)\ \ \ \text{for final state,}$,
[…]
$k^{M}\mathcal{M}_{M}=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for incoming,}\qquad k^{*M}\mathcal{M}_{M}=0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{for outgoing,}$, (6)
with $\mathcal{M}^{M}\equiv(\mathcal{M}^{\mu},\mathcal{M}(\varphi))$ and the 5-component “momentum” $k^{M}\equiv(k^{\mu},-im_{V})$. The 5-dimensional “metric” for index contraction is $g_{MN}\equiv\operatorname{diag}(+1,-1,-1,-1,-1)$. The spin sum of the polarization vectors becomes
@@
$\epsilon_{\mathrm{L}}^{M}(k,\lambda)\equiv\left(-\frac{m_{V}n^{\mu}}{n\cdot k},\leavevmode\nobreak\ i\right)+\lambda\,\frac{k^{M}}{m_{V}}$ (9)
gives the same amplitudes for any value of $\lambda$. Choosing $\lambda=1$ yields $\epsilon_{\mathrm{L}}^{M}(k,1)=(\epsilon_{\mathrm{L}}^{\mu}(k),0)$, which corresponds to the conventional case using Eq. (3). We will refer to the case of $\lambda=1$ as the gauge form of the longitudinal polarization. The case of $\lambda=0$ corresponds to Eq. (5), which we will call the GE form.
@@ II.2 Gauge choice and Propagator
The 5-component formalism can also be naturally derived from the Lagrangian, which includes a term $m_{V}V_{\mu}\partial^{\mu}\varphi$ that mixes the gauge field with the Goldstone field after gauge symmetry breaking Chen:2022gxv . Therefore, it is natural to combine gauge and Goldstone components into one 5-component object.
The GE representation, implemented through the 5-component formalism, is compatible with any gauge except the unitary gauge, which eliminates Goldstone bosons from the Feynman rules. In the $R_{\xi}$ gauges, the gauge-fixing Lagrangian term $-(\partial^{\mu}V_{\mu}-m_{V}\varphi)^{2}/(2\xi)$ is introduced to cancel the gauge-Goldstone mixing term. This results in the propagator for a massive vector boson in the 5-component formalism taking the form of
@@
To completely eliminate the $k^{\mu}$ terms, we can choose the gauge condition $n^{\mu}(x)V_{\mu}(x)=0$, where $n^{\mu}(x)$ is the Fourier transformation of $n^{\mu}(k)=(1,-{\vec{k}}/{|\vec{k}|})$, by adding the Lagrangian term $-[n^{\mu}(x)V_{\mu}(x)]^{2}/(2\alpha)$. Its specific form is irrelevant here. Applying to a massive gauge theory, the 5-dimensional vector boson propagator becomes
@@ II.3 Vertices and Feynman Rules
With the polarization vectors and the propagator expressed in the 5-component formalism, it is natural to extend this framework to vertices involving massive vector bosons. To reflect this property, we introduce a double-line notation by overlapping a wavy line with a dashed line to represent the massive vector boson in Feynman rules: the wavy line represents the gauge components, while the dashed line represents the Goldstone component. As examples of this double-line notation, we illustrate the decomposition of the vector boson propagator in Fig. 1 and that of the $WWh$ vertex in Fig. 2.
Figure 1: 5-component vector boson propagator in the double-line notation.
Figure 2: 5-component vertex of $WWh$ in the double-line notation. The minus sign in front of $\varphi Wh$ comes from the $g_{44}=-1$ component of the “metric” $g_{MN}$.
The new form of vertices brings additional subtleties. As an example, we denote the $WWh$ vertex to be $V_{WWh}^{MN}$, where the subscript indicates the vertex and the superscript represents indices. In the GE representation, a vertex involving massive vector bosons has multiple parameters. For instance, $V^{MN}_{WWh}$ not only includes the pure gauge components $V^{\mu\nu}_{WWh}$ with coupling $g_{WWh}$, but also involves the gauge-Goldstone components $V^{\mu 4}_{WWh}=V^{\mu}_{W\varphi h}$ with coupling $g_{W\varphi h}=g_{\varphi Wh}$ and the pure Goldstone component $V_{WWh}^{44}=V_{\varphi\varphi h}$ with coupling $\lambda_{\varphi\varphi h}$. We can further express the vertex as $V^{MN}_{WWh}(g_{WWh},g_{\varphi Wh},\lambda_{\varphi\varphi h};m_{h},m_{W})$. These couplings, however, are not independent in the SM. They are all determined by the gauge coupling $g$, excluding the mass parameters. These relations come from the spontaneous symmetry breaking of the EW theory, reflecting of the underlying gauge symmetry. Similar relations also exist for other vertices such as $ff^{\prime}W$, $WWZ$, etc. Further details will be discussed in Sec. IV.
Finally, let us examine the free parameters in the SM. The EW gauge and Higgs sectors of the SM contain only four free parameters, which we choose to be $m_{W}$, $g$, $\theta_{\mathrm{W}}$, and $m_{h}$, representing the $W$ boson mass, the $\mathrm{SU}(2)_{\mathrm{L}}$ gauge coupling, the weak mixing angle, and the Higgs boson mass, respectively. All other parameters in these sectors are derived from them. For example, the $Z$ boson mass is given by $m_{Z}=m_{W}/\cos\theta_{\mathrm{W}}$, and the Higgs self-coupling is $\lambda_{h}=2m_{h}/v^{2}=g^{2}m_{h}/(2m_{W}^{2})$. In the fermion sector, each Yukawa coupling is defined as $\lambda_{f}=gm_{f}/(\sqrt{2}m_{W})$, introducing an additional parameter, the corresponding fermion mass $m_{f}$. If we assume that the particle masses have been precisely measured and are treated as inputs, then the free parameters in the EW sectors of the SM reduce to just two: $g$ and $\theta_{\mathrm{W}}$.

### v3

# Numerical study on the gauge symmetry of electroweak amplitudes

@@ Abstract
Electroweak (EW) amplitudes in the gauge-Goldstone five-component formalism have a distinctive property: gauge symmetry is imprinted in the amplitudes, manifested as the massive Ward identity (MWI) $k^{M}\mathcal{M}_{M}=0$. In this study, we used the HELAS package to numerically study gauge symmetry in EW amplitudes. First, we directly tested gauge symmetry by examining the MWI of amplitudes. Second, we modified the couplings within a vertex and among vertices to check if and how the MWI changes. Third, we tested gauge symmetry by considering the couplings modified by operators from the standard model effective field theory (SMEFT). Similar to the standard model, there are relations between different couplings that are protected by gauge symmetry. We observed that, if we modify the couplings to deviate from the relations, the MWI is violated. In contrast, the MWI is restored when the relations between couplings reduce to those in the SMEFT.
@@ I Introduction
The gauge cancellation of electroweak (EW) scattering amplitudes in the standard model (SM) Glashow:1961tr ; Weinberg:1967tq ; Salam:1968rm has been both a theoretical problem hindering physical analysis and a practical problem in numerical calculations. In recent years, a new framework for computing EW amplitudes based on Goldstone equivalence has been proposed and implemented in the software package HELAS Chen:2022gxv . The main ingredients of this framework are as follows: 1) taking Goldstone equivalence Cornwall:1974km ; Chanowitz:1985hj ; Gounaris:1986cr , so that the polarization vector of a massive vector boson has no $k^{\mu}$ term and is instead composed of its Goldstone component and a remnant gauge term; 2) combining the gauge components and their Goldstone components of fields or polarization vectors into single five-component objects at the level of polarization vectors, propagators, and vertices; and 3) imposing a special light-cone gauge defined by the gauge direction $n^{\mu}=(1,-{\vec{k}}/{|\vec{k}|})$, dubbed the Feynman diagram gauge Chen:2022gxv or Goldstone equivalence gauge Chen:2016wkt . It has been demonstrated in Ref. Chen:2022gxv with many examples that EW amplitudes in this new scheme do not have gauge cancellation, thereby solving a long-standing problem. Further studies on this topic can be found in Refs. Hagiwara:2024ron ; Chen:2022xlg ; Furusato:2024uli ; Hagiwara:2024xdh ; Zheng:2024lwl ; Furusato:2024uli ; Furusato:2024ghr ; Jeong:2025hwj . Earlier but incomplete treatment of the five-component formalism can be found in Kunszt:1987tk ; Wulzer:2013mza ; Cuomo:2019siu .
@@
The new framework in Ref. Chen:2022gxv is essentially a reorganization and combination of “pieces” from existing Feynman rules. Although ideal to be combined with the Feynman diagram gauge, it can also be applied to other gauges, such as the Feynman gauge. Therefore, it is necessary to distinguish the five-component framework from any specific choice of gauge. Additionally, since the framework relies on the Goldstone equivalence theorem Cornwall:1974km ; Chanowitz:1985hj ; Gounaris:1986cr in an essential way, it is appropriate to refer to it as the Goldstone equivalence (GE) representation of Feynman rules, as we will do from now on. Correspondingly, the standard Feynman rules will be called the gauge representation, as they exclusively treat the physical content of massive vector bosons as quanta of gauge fields.
@@
Although the absence of gauge cancellation is crucial for practical applications, it is not the only important property of the GE representation of EW Feynman rules. Another intriguing property is that the GE representation directly imprints gauge symmetry at the level of amplitudes, as expressed by the massive Ward identity (MWI), which is named in analogue to its massless counterpart. Specifically, for a scattering amplitude $\varepsilon^{\mu(*)}(k)\mathcal{M}_{\mu}$ involving an external vector boson $V$ with mass $m_{V}$ and four-momentum $k^{\mu}$, the MWI is Cornwall:1974km ; Chanowitz:1985hj ; Gounaris:1986cr ; Bagger:1989fc
@@
where the - (++) sign corresponds to the case where the vector boson is in the initial (final) state, and $\mathcal{M}(\varphi)$ represents the amplitude obtained by replacing the vector boson with the corresponding Goldstone boson $\varphi$. The MWI (1) is a fundamental identity for a spontaneously broken gauge theory, e.g., the EW gauge theory. This study aimed to investigate the gauge symmetry of EW amplitudes using the numerical tool HELAS Hagiwara:1990dw ; Murayama:1992gi ; Chen:2022gxv .
Similar to the Ward identity in gauge theories with massless gauge bosons, the MWI guarantees precise relations among different diagrams of an amplitude, which further imply exact relations among different vertices or, equivalently, different couplings within the same theory. Considering the example of $W^{+}W^{-}\rightarrow W^{+}W^{-}$ with the helicity combination of $\mathrm{TTTT}$, where $\mathrm{T}$ represents a transverse polarization, the amplitude includes three vertices $WWZ$, WWAWWA, and $WWWW$ with couplings $g_{WWZ}$, $g_{WWA}$, and $g_{WWWW}$. The Ward identity requires that
@@
Conversely, any deviation from this relation would result in violations of the gauge symmetry and the Ward identity. When the helicity combination involves longitudinal components, the precise relations the MWI brings about not only involve couplings of gauge bosons, but also those of Goldstone bosons and the Higgs boson. Thus, the structure of gauge symmetry is much richer in massive amplitudes, which we study in detail. Moreover, the gauge symmetry and MWI are not limited to the SM and can be applied to other theories. One theory of particular interest is the standard model effective field theory (SMEFT) Coleman:1969sm ; Callan:1969sn ; Weinberg:1980wa ; Leung:1984ni ; Buchmuller:1985jz ; Grzadkowski:2010es , which is a direct extension of the SM.
The remainder of this paper is organized as follows. In Section II, we provide a brief introduction to the GE representation of EW interactions, emphasizing the relations among different couplings and parameters. In Section III, we directly test the EW gauge symmetry with the MWI (1) across multiple processes and helicity combinations. In Section IV, we examine the EW gauge symmetry through anomalous couplings, including both within individual vertices and among multiple vertices. Finally, in Section V, we study the connection between anomalous couplings (Higgs self-couplings and Yukawa couplings) and certain SMEFT operators using the MWI. We demonstrate that, when the anomalous couplings are adjusted according to the SMEFT operators, the MWI is satisfied, ensuring the gauge invariance and self-consistency of the theory.
@@ II GE representation of EW interactions
II.1 Polarization Vectors in the Five-Component Formalism
The central identity of our program is the MWI (1), which essentially states the equivalence between the $k^{\mu}$ terms in gauge fields and the corresponding Goldstone fields. Moreover, the longitudinal polarization vector of a massive vector boson $V$ with four-momentum $k^{\mu}$ can be decomposed into
@@ II.1 Polarization Vectors in the Five-Component Formalism
with $n^{\mu}=(1,-{\vec{k}}/{|\vec{k}|})$, and the subscript $\mathrm{L}$ representing a longitudinal polarization. We can eliminate the $k^{\mu}$ term in $\epsilon_{\mathrm{L}}^{\mu}(k)$ by using the MWI, resulting in the amplitude with a longitudinal vector boson:
@@
The generalization to multiple vector bosons is straightforward, but it leads to an excessive number of terms, making calculations cumbersome. To address this issue, we define five-component longitudinal polarization vectors in the GE representation by combining the gauge boson wave function and the Goldstone boson “wave function” as
$\epsilon^{M}_{\mathrm{L}}(k)$ $\equiv\left(-\frac{m_{V}n^{\mu}}{n\cdot k},\penalty 10000\ i\right)\ \ \ \text{for initial state,}$ (5)
$\epsilon^{*M}_{\mathrm{L}}(k)$ $\equiv\left(-\frac{m_{V}n^{\mu}}{n\cdot k},\penalty 10000\ -i\right)\ \ \ \text{for final state,}$
with $\mu=0,1,2,3$ and $M=0,1,2,3,4$. In addition, the five-component transverse polarization vectors are defined as $\epsilon^{M}_{\pm}(k)=\big(\epsilon^{\mu}_{\pm}(k),\penalty 10000\ 0\big)$. Then, the MWI (1) can be rewritten as
$k^{M}\mathcal{M}_{M}=0\penalty 10000\ \penalty 10000\ \text{for incoming,}\qquad k^{*M}\mathcal{M}_{M}=0\penalty 10000\ \penalty 10000\ \text{for outgoing,}$ (6)
with $\mathcal{M}^{M}\equiv(\mathcal{M}^{\mu},\mathcal{M}(\varphi))$, and the five-component “momentum” $k^{M}\equiv(k^{\mu},-im_{V})$. The five-dimensional “metric” for index contraction is $g_{MN}\equiv\operatorname{diag}(+1,-1,-1,-1,-1)$. The spin sum of the polarization vectors becomes
@@
$\epsilon_{\mathrm{L}}^{M}(k,\lambda)\equiv\left(-\frac{m_{V}n^{\mu}}{n\cdot k},\penalty 10000\ i\right)+\lambda\,\frac{k^{M}}{m_{V}}$ (9)
which gives the same amplitudes for any value of $\lambda$. Choosing $\lambda=1$ yields $\epsilon_{\mathrm{L}}^{M}(k,1)=(\epsilon_{\mathrm{L}}^{\mu}(k),0)$, which corresponds to the conventional case using Eq. (3). We refer to the case of $\lambda=1$ as the gauge form of the longitudinal polarization. The case of $\lambda=0$ corresponds to Eq. (5), which we call the GE form.
@@ II.2 Gauge Choice and Propagator
The five-component formalism can also be naturally derived from the Lagrangian, which includes a term $m_{V}V_{\mu}\partial^{\mu}\varphi$ that mixes the gauge field with the Goldstone field after gauge symmetry breaking Chen:2022gxv . Therefore, it is natural to combine gauge and Goldstone components into one five-component object.
The GE representation, implemented through the five-component formalism, is compatible with any gauge except the unitary gauge, which eliminates Goldstone bosons from the Feynman rules. In the $R_{\xi}$ gauges, the gauge-fixing Lagrangian term $(\partial^{\mu}V_{\mu}-m_{V}\varphi)^{2}/(2\xi)$ is introduced to cancel the gauge-Goldstone mixing term. This results in the propagator for a massive vector boson in the five-component formalism taking the form of
@@
To completely eliminate the $k^{\mu}$ terms, we can choose the gauge condition $n^{\mu}(x)V_{\mu}(x)=0$, where $n^{\mu}(x)$ is the Fourier transformation of $n^{\mu}(k)=(1,-{\vec{k}}/{|\vec{k}|})$, by adding the Lagrangian term $n^{\mu}(x)V_{\mu}(x)]^{2}/(2\alpha)$. Its specific form is irrelevant here. Applying this to a massive gauge theory, the five-dimensional vector boson propagator becomes
@@ II.3 Vertices and Feynman Rules
With the polarization vectors and propagator expressed in the five-component formalism, this framework can be naturally extended to vertices involving massive vector bosons. Accordingly, we introduce a double-line notation by overlapping a wavy line with a dashed line to represent the massive vector boson in Feynman rules: the wavy line represents the gauge components, whereas the dashed line represents the Goldstone component. As examples of this double-line notation, we illustrate the decomposition of the vector boson propagator in Fig. 1 and that of the $WWh$ vertex in Fig. 2.
Figure 1: Five-component vector boson propagator in the double-line notation.
Figure 2: Five-component vertex of $WWh$ in the double-line notation. The minus sign before $\varphi Wh$ comes from the $g_{44}=-1$ component of the “metric” $g_{MN}$.
The new form of vertices brings additional subtleties. For example, we denote the $WWh$ vertex as $V_{WWh}^{MN}$, where the subscript indicates the vertex and the superscript represents indices. In the GE representation, a vertex involving massive vector bosons has multiple parameters. For instance, $V^{MN}_{WWh}$ not only includes the pure gauge components $V^{\mu\nu}_{WWh}$ with coupling $g_{WWh}$ but also involves the gauge-Goldstone components $V^{\mu 4}_{WWh}=V^{\mu}_{W\varphi h}$ with coupling $g_{W\varphi h}=g_{\varphi Wh}$ and the pure Goldstone component $V_{WWh}^{44}=V_{\varphi\varphi h}$ with coupling $\lambda_{\varphi\varphi h}$. We can further express the vertex as $V^{MN}_{WWh}(g_{WWh},g_{\varphi Wh},\lambda_{\varphi\varphi h};m_{h},m_{W})$. However, these couplings are not independent in the SM. They are all determined by the gauge coupling $g$, excluding the mass parameters. These relations originate from the spontaneous symmetry breaking of the EW theory, reflecting of the underlying gauge symmetry. Similar relations also exist for other vertices such as $ff^{\prime}W$ and $WWZ$. Further details are discussed in Sec. IV.
Finally, let us examine the free parameters in the SM. The EW gauge and Higgs sectors of the SM contain only four free parameters, which we choose to be $m_{W}$, $g$, $\theta_{\mathrm{W}}$, and $m_{h}$, representing the $W$ boson mass, $\mathrm{SU}(2)_{\mathrm{L}}$ gauge coupling, weak mixing angle, and Higgs boson mass, respectively. All the other parameters in these sectors are derived from them. For example, the $Z$ boson mass is given by $m_{Z}=m_{W}/\cos\theta_{\mathrm{W}}$, and the Higgs self-coupling is $\lambda_{h}=2m_{h}/v^{2}=g^{2}m_{h}/(2m_{W}^{2})$. In the fermion sector, each Yukawa coupling is defined as $\lambda_{f}=gm_{f}/(\sqrt{2}m_{W})$, introducing an additional parameter, the corresponding fermion mass $m_{f}$. If we assume that the particle masses have been precisely measured and are treated as inputs, the free parameters in the EW sectors of the SM reduce to just two: $g$ and $\theta_{\mathrm{W}}$.